OpenStudy (anonymous):

you make 60 ounces of dough for breadsticks. each breadstick requires 5/4 ounces of dough. how many breadsticks can u make

OpenStudy (ddcamp):

How many times does 5/4 go into 60?

OpenStudy (anonymous):

0.025

OpenStudy (anonymous):

\[\frac{60}{\frac{5}{4}}=48 \]

OpenStudy (anonymous):

i have to turn it in turn it in to a fraction and show my work. i hate stupid word problems

OpenStudy (ddcamp):

Turning it into a fraction isn't too bad: \[60 \rightarrow \frac{60}{1}\] Then we have: \[\frac{60}{1} \div \frac{5}{4} \rightarrow \frac{60}{1} \times \frac{4}{5}\]

OpenStudy (anonymous):

i cross-multiplied and still wound up with the wrong answer

OpenStudy (anonymous):

never mind thanks guys, i was lookin at the wrong answer in the back of the book